dy/dx = 5y + (e ^−2x)(y^−2)


1
Expert's answer
2022-02-18T12:38:32-0500

Solution:

"\begin{aligned}

&\frac{d y}{d x}-5 y=e^{-2 x} y^{-2} \\

&\Longrightarrow y^{2} \frac{d y}{d x}-5 y^{3}=e^{-2 x} \\

&\Longrightarrow 3 y^{2} \frac{d y}{d x}-15 y^{3}=3 e^{-2 x} \\

&\text { Let, } \mathrm{z}=\mathrm{y}^{\wedge} 3 \\

&\frac{d z}{d x}=3 y^{2} \\

&\Longrightarrow \frac{d z}{d x}-15 z=3 e^{-2 x} \ldots(i)

\end{aligned}"

Now this differential equation is of the form

"\frac{d y}{d x}+P(x) y=Q(x)"

here, "P(x)=-15"

"Q(x)=3 e^{-2 x}"

So,

"\text { Integrating Factor }=e^{\int P(x) d x}=e^{\int-15 d x}=e^{-15 x}"

Multiplying Integrating factor to the both sides of equation (i),

 "e^{-15 x} \frac{d z}{d x}-15 e^{-15 x} z=3 e^{-17 x}"

"\begin{aligned}

&\Longrightarrow \frac{d\left(e^{-15 x} z\right)}{d x}=3 e^{-17 x} \\

&\Longrightarrow d\left(e^{-15 x} z\right)=3 e^{-17 x} d x

\end{aligned}"

Integrating both sides,

"\begin{aligned}

&\Longrightarrow \int d\left(e^{-15 x} z\right)=\int 3 e^{-17 x} d x \\

&\Longrightarrow e^{-15 x} z=-\frac{3}{17} e^{-17 x}+C \\

&\Longrightarrow y^{3} e^{-15 x}=-\frac{3}{17} e^{-17 x}+C 

\end{aligned}"